I guessed on this one. 1- 0.843 = 0.157 Is this correct or how do you work this problem out. Thanks Business Analysis 2372 homework.

Netflix Inc. is a large and extraordinarily popular streaming media and video-on-demand provider accessible to customers in almost every country in the world. Part of the reason for its success has been its adoption of business analytics, the use of statistical analysis techniques to extract information from data that can be used in strategic planning.

Key components of the Netflix supply chain are the "content partners," studios or production companies that provide the audio and video for media as well as supporting content such as subtitles and closed captioning (the company calls these elements, collectively, "assets.")  A recent issue facing the company was inconsistent quality in the assets it was receiving from some of its content partners. Using a technique called "predictive quality control," Netflix was able to build a statistical model that could predict the probability of a defect in an asset based on several factors, such as the age and the genre of the media it was associated with.

Let D be the event {an asset is defective}, and suppose the model predicts that P(D) = 0.843. What is P(Dc), the complement of the event D, to three decimal places?
Answer #1

From probability axiom we know that

P(D) = 1-P(D) = 1-0.843 0.157